OpenStudy (anonymous):

Write the explicit formula that represents the geometric sequence -2, 8, -32, 128. what am i supposed to do with the negative numbers? im so confused, please help

myininaya (myininaya):

what is the common ratio?

myininaya (myininaya):

just take a term and divide it by it's previous term to find

OpenStudy (anonymous):

would it be 4?

myininaya (myininaya):

well 8/-2 is -4 actually

myininaya (myininaya):

\[a_n=ar^{n-1}\]

myininaya (myininaya):

replace r with -4

OpenStudy (anonymous):

so is that how you found the common ratio?

myininaya (myininaya):

yep just take a term and divide it by it's previous term

OpenStudy (anonymous):

xn = 8 × -2(n-1)?

OpenStudy (anonymous):

8 x -4

myininaya (myininaya):

\[a_n=a \cdot r^{n-1 } \\ r \text{ is common ratio number } \\ a \text{ is first term in the sequence }\]

OpenStudy (anonymous):

so is that all you have to do? or do you ri==write out an equation for it?

myininaya (myininaya):

\[a_n=a \cdot r^{n-1 } \\ r \text{ is common ratio number } \\ a \text{ is first term in the sequence }\] all you do is replace r with -4 and a with the first number in the sequence

myininaya (myininaya):

do you see what your first number is in your sequence?

OpenStudy (anonymous):

okay thanks!

myininaya (myininaya):

we have a=-2 and r=-4 so our formula is: \[a_n=(-2) \cdot (-4)^{n-1} \]